1.the angle of elevation of the top of a flag post from a point on the ground level 40m away from the flag post is 60 degree. find the hieght of the post
Found 2 solutions by fawadhaider, xcentaur:
Answer by fawadhaider(1)

As we know,
Tan <=P/B
putting the values
Tan60=P/40
P=40Tan60
Height of flag post= 40tan60 Answer

Answer by xcentaur(357)

Let the height be h m.
Then,
Tan60=h/40
Height of flag post= 40(tan60)
=40(sqrt3)
=40(1.732)
=69.28m
=70m


Prabhat.